# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 4.2. Local-energy-preserving approximation
(4.16) $\mu^{\sigma}(K_{\varepsilon})\geq 1-\varepsilon,\text{ for }0<\sigma<1.$
@@
$K_{R}$ :=$:=\Big{\{}({{\bf{v}}},{\bf{d}})\Big{|}\|{{\bf{v}}}\|_{L^{\infty}(0,T;H)\cap L^{2}(0,T;V)}+\|\partial_{t}{{\bf{v}}}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
@@
By the Aubin-Lions lemma, we conclude that $K_{R}$ is compact in $L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})$ with $1<p_{1}<\frac{10}{3}$, and $10\displaystyle 1<p_{2}<$ 10. Moreover, by the estimates (3.5), (4.11), (4.14) and (4.15), we can follow the same argument as in [19, (381)-(383)] to conclude that there exists a sufficiently large $R>0$ such that $K_{1/R}$ satisfies (4.16). Consequently, the laws of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$, which will be denoted as $\xi^{\sigma}$ subsequentely, are tight over the phase space $\mathcal{G}$:
@@
Since the family of measures $\{\xi^{\sigma}\}$ is weakly compact on $\mathcal{G}$, we deduce that $\{\xi^{\sigma}\}$ converges weakly to a probability measure $\xi$ on $\mathcal{G}$ up to a subsequence. Hence, using a version of the Skorokhod embedding theorem (See [42, Lemma 3.3]), we can conclude that there exists a probability tripe $(\widetilde{\Omega},\widetilde{\mathcal{F}},\widetilde{\mathbb{P}})$, a sequence of random variables $\widetilde{U}_{k}=(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{\pi}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k},\widetilde{W}^{k})$ with values in $\mathcal{E}$, and an element $\widetilde{U}=(\widetilde{{\bf{v}}},\widetilde{\pi},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})\in\mathcal{E}$ such that for each $k$, $\widetilde{U}_{k}$ has the same law with $U_{\sigma_{k}}$, in particular, $\widetilde{W}^{k}$ is a Brownian motion adapted to the filtration {~$\{\widetilde{\mathcal{F}}^{k}_{t}\}_{t\geq 0}$ which is given by the completion of the $\sigma$-algebra generated by $\{\widetilde{U}_{k}(s);s\leq t\}$, and
(4.17) $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k},\widetilde{W}^{k})\rightarrow(\widetilde{{\bf{v}}},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})\quad\widetilde{\mathbb{P}}-a.s.,\mbox{ as }k\to\infty,$
@@
in the topology of $\mathcal{G}$. Additionally, setting {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ as the completion of the $\sigma$-algebra generated by $\{\widetilde{U}(s);s\leq t\}$, we obtain a new stochastic basis $\widetilde{\mathfrak{B}}=\left(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t)\right)_{t\geq 0})$, as in (3.7).
@@
Next we show that for each $k$, the random variable $(\widetilde{{\bf{v}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{\bf{d}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfies the local energy equality (4.1). We adapt a trick of Bensoussan [2]. Given $\varphi\in C_{c}^{\infty}(Q_{T};\mathbb{R}_{+})$, for $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{{\bf{z}}}})$, the components of ${U}_{\sigma}$, define the random variable $X^{\sigma}$ on $(\Omega,\mathcal{F})$ as
@@
$\displaystyle X^{\sigma}$ $\displaystyle=\Bigg{|}\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{Q_{T}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
@@
$\displaystyle-\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
@@
$\displaystyle-\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi-{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi-\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
@@
$\displaystyle-\int_{Q_{T}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle-\int_{Q_{T}}({{\bf{z}}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))-\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle+\int_{Q_{T}}f({{\bf{d}}}^{\sigma})(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})+2\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi\Bigg{|}.$
Following a similar proof as for Lemma 4.1, we can conclude that for $1\displaystyle 0<\sigma<$ 1, we have
$\displaystyle\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{\mathbb{T}^{3}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
$\displaystyle\leq\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi+{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle+\int_{\mathbb{T}^{3}}({{\bf{z}}}\cdot{\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])})\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))+[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle-\int_{\mathbb{T}^{3}}f({{\bf{d}}}^{\sigma})\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})-2\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
Hence $X^{\sigma}=0$ with probability 1. Thus we have $\mathbb{E}(g(U_{\sigma}))=0$, where $g(U_{\sigma}):=X^{\sigma}/(1+X^{\sigma}).$ Let $\widetilde{X}^{k}$ be defined analogously for $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$, the components of $\widetilde{U}_{k}$, as a random variable on $(\widetilde{\Omega},\widetilde{\mathcal{F}})$. Note that $g(\cdot)$ is a deterministic bounded continuous function on a subset of $\mathcal{E}$, and thus
@@
as desired. With a similar method, we can verify one more thing. That is, for each $\sigma_{k}$, $\widetilde{U}_{k}$ satisfies the equations (3.2) and (4.9) in the sense of distribution $\widetilde{\mathbb{P}}$-a.s.. Particularly, the random variables $(\widetilde{{{\bf{v}}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{{\bf{d}}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfy the approximation equation (4.9) $\widetilde{\mathbb{P}}$-a.s., that is
@@
(4.19) $\left\{\begin{array}[]{l}\partial_{t}\widetilde{{\bf{v}}}^{\sigma_{k}}-\Delta\widetilde{{\bf{v}}}^{\sigma_{k}}+(\widetilde{{\bf{z}}}^{k}+\Phi_{\sigma_{k}}[\widetilde{{\bf{v}}}^{\sigma_{k}}])\cdot\nabla(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})+\nabla\widetilde{\pi}^{\sigma_{k}}=-\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]\cdot(\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}})),\\ \nabla\cdot\widetilde{{\bf{v}}}^{\sigma_{k}}=0,\\ \partial_{t}\widetilde{{\bf{d}}}^{\sigma_{k}}-(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})\cdot\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]=\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}}).\end{array}\right.$
As a result, passing to the limit on (the weak form of) the equations (3.2) and (4.19) as $\sigma_{k}\to 0$, we conclude that $(\widetilde{{\bf{v}}},\widetilde{{\bf{d}}},\widetilde{{\bf{z}}},\widetilde{W})$ solves (3.2) and (3.3) in the sense of distribution (see Def. 3.1) $\widetilde{\mathbb{P}}$-a.s.. Thus there exists a distribution $\widetilde{\pi}$ defined on the stochastic basis $\widetilde{\mathfrak{B}}$ which plays the role of the pressure field for $\widetilde{{\bf{v}}}$. In fact, let $\widetilde{{\pi}}$ solve the following Poisson equation
$-\Delta\widetilde{{\pi}}=\mathop{\rm div}\nolimits^{2}[(\widetilde{{{\bf{z}}}}+\widetilde{{{\bf{v}}}})\otimes(\widetilde{{{\bf{z}}}}+\widetilde{{{\bf{v}}}})]+\mathop{\rm div}\nolimits(\nabla\widetilde{{\bf{d}}}\cdot(\Delta\widetilde{{\bf{d}}}-f(\widetilde{{\bf{d}}}))),$
with $\int_{{\mathbb{T}}^{3}}\widetilde{\pi}(t)=0$ for every $t$. Thanks to (4.19), for almost sure $\widetilde{\omega}\in\widetilde{\Omega}$, we have the same estimate as in (4.13) for $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})$, which implies that the sequence $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})$ is bounded uniformly in $L^{5/3}(Q_{T})$ independent of $\sigma$. Hence we have for almost every fixed $\widetilde{\omega}\in\widetilde{\Omega}$, there exists a subsequence that may depend on $\widetilde{\omega}$ such that
(4.20) $\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})\to\widetilde{\pi}(\widetilde{\omega})\mbox{ weakly in }L^{\frac{5}{3}}(Q_{T})\mbox{ up to the subsequence}.$
Finally, we show that $(\widetilde{{\bf{v}}},\widetilde{\pi},\widetilde{{\bf{d}}})$ satisfies the local energy inequality described in Definition 3.2. For almost sure $\widetilde{\omega}\in\widetilde{\Omega}$, as proven above (see equation (4.18)) we know that the random variables $(\widetilde{{\bf{v}}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{\bf{d}}}^{\sigma_{k}}(\widetilde{\omega}),\widetilde{{{\bf{z}}}}^{k}(\widetilde{\omega}))$ satisfy the local energy equality (4.1). Thus for each such $\widetilde{\omega}$, passing to the limit upon the corresponding subsequence as in (4.20)(using also (4.17)), we obtain the local energy inequality for $(\widetilde{{\bf{v}}}(\widetilde{\omega}),\widetilde{\pi}(\widetilde{\omega}),\widetilde{{\bf{d}}}(\widetilde{\omega}),\widetilde{{{\bf{z}}}}(\widetilde{\omega}))$ by the lower semicontinuity.
5. Partial regularity
5.1. Blowing up analysis

### v2

@@ 4.2.1 Compactness of the approximation sequence
$\mu^{\sigma}(K_{\varepsilon})\geq 1-\varepsilon,\text{ for }0<\sigma<1.$ (38)
@@
$K_{R}$ $=\Big\{({{\bf{v}}},{\bf{d}})\Big|\|{{\bf{v}}}\|_{L^{\infty}(0,T;H)\cap L^{2}(0,T;V)}+\|\partial_{t}{{\bf{v}}}\|_{L^{2}(0,T;D(A^{-1}))+L^{2}(0,T;H^{-1}(\mathbb{T}^{3}))+L^{\frac{5}{4}}(Q_{T})}$
@@
By the Aubin-Lions lemma, we conclude that $K_{R}$ is compact in $L^{p_{1}}(Q_{T})\times L^{p_{2}}(Q_{T})$ with $1<p_{1}<\frac{10}{3}$, and $1<p_{2}<10$. Moreover, by the estimates (78), (25), (36) and (37), we can follow the same argument as in [25, (381)-(383)] to conclude that there exists a sufficiently large $R>0$ such that $K_{1/R}$ satisfies (38). Consequently, the laws of $({{\bf{v}}}^{\sigma},{{\bf{d}}}^{\sigma},{{\bf{z}}},W)$, which will be denoted as $\xi^{\sigma}$ subsequently, are tight over the phase space $\mathcal{G}$:
@@
[…]
@@
in the topology of $\mathcal{G}$. Additionally, setting $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ as the completion of the $\sigma$-algebra generated by $\{\widetilde{U}(s);s\leq t\}$, we obtain a new stochastic basis $\widetilde{\mathfrak{B}}=\left(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},\left(\widetilde{W}(t)\right)_{t\geq 0}\right)$, as in (16).
4.2.2 Preservation of the key energy equality
@@
Next we show that for each $k$, the random variable $(\widetilde{{\bf{v}}}^{\sigma_{k}},\widetilde{{\pi}}^{\sigma_{k}},\widetilde{{\bf{d}}}^{\sigma_{k}},\widetilde{{{\bf{z}}}}^{k})$ satisfies the local energy equality (81). We adapt a technique due to Bensoussan [4]. Given $\varphi\in C_{0}^{\infty}(Q_{T};\mathbb{R}_{+})$, for $({{\bf{v}}}^{\sigma},{\pi}^{\sigma},{{\bf{d}}}^{\sigma},{{{\bf{z}}}})$, the components of ${U}_{\sigma}$, define the random variable $X^{\sigma}$ on $(\Omega,\mathcal{F})$ as
$\displaystyle X^{\sigma}$ $\displaystyle:=\left|X_{1}^{\sigma}-X_{2}^{\sigma}\right|,$ (40)
where
@@
$\displaystyle X_{1}^{\sigma}$ $\displaystyle:=\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\varphi+\int_{Q_{T}}(|\nabla{{\bf{v}}}^{\sigma}|^{2}+|\Delta{{\bf{d}}}^{\sigma}|^{2}+|f({{\bf{d}}}^{\sigma})|^{2})\varphi$
and
@@
$\displaystyle X_{2}^{\sigma}$ $\displaystyle:=\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}+F({{\bf{d}}}^{\sigma})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla{{\bf{v}}}^{\sigma}]\cdot{{\bf{z}}}\varphi$
@@
$\displaystyle+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+{{\bf{v}}}^{\sigma}\cdot{\bf{z}}\right)({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla\varphi+{\pi}^{\sigma}{{\bf{v}}}^{\sigma}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}^{\sigma}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}^{\sigma}|^{2}}{2}\right)\Delta\varphi$
@@
$\displaystyle+\int_{Q_{T}}\left(\nabla{{\bf{d}}}^{\sigma}\odot\nabla{{\bf{d}}}^{\sigma}-\frac{1}{2}|\nabla{{\bf{d}}}^{\sigma}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$\displaystyle+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}])\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}))+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla{{\bf{d}}}^{\sigma}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}^{\sigma}$
$\displaystyle-\int_{Q_{T}}f({{\bf{d}}}^{\sigma})(\nabla\varphi\cdot\nabla{{\bf{d}}}^{\sigma})-2\nabla f({{\bf{d}}}^{\sigma}):\nabla{{\bf{d}}}^{\sigma}\varphi.$
From (81) in Lemma 13, we infer that
$\displaystyle X_{1}^{\sigma}=X_{2}^{\sigma}\mbox{ with Probability 1}.$ (41)
[…]
@@
as required.
4.2.3 Passage to the limit
[…]
@@ 4.2.3 Passage to the limit
$\begin{cases}&\partial_{t}\widetilde{{\bf{v}}}^{\sigma_{k}}-\Delta\widetilde{{\bf{v}}}^{\sigma_{k}}+(\widetilde{{\bf{z}}}^{k}+\Phi_{\sigma_{k}}[\widetilde{{\bf{v}}}^{\sigma_{k}}])\cdot\nabla(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})+\nabla\widetilde{\pi}^{\sigma_{k}}\\ &\quad\quad\quad\quad\quad\quad=-\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]\cdot(\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}})),\\ &\nabla\cdot\widetilde{{\bf{v}}}^{\sigma_{k}}=0,\\ &\partial_{t}\widetilde{{\bf{d}}}^{\sigma_{k}}-(\widetilde{{\bf{z}}}^{k}+\widetilde{{\bf{v}}}^{\sigma_{k}})\cdot\nabla\Phi_{\sigma_{k}}[\widetilde{{\bf{d}}}^{\sigma_{k}}]=\Delta\widetilde{{\bf{d}}}^{\sigma_{k}}-f(\widetilde{{\bf{d}}}^{\sigma_{k}}).\end{cases}$ (43)
[…]
$\widetilde{{\pi}}^{\sigma_{k}}(\widetilde{\omega})\to\widetilde{\pi}(\widetilde{\omega})\mbox{ weakly in }L^{\frac{5}{3}}(Q_{T})\mbox{ up to the subsequence}.$ (44)
[…]
5 Partial regularity of the martingale suitable weak solutions
With the global existence of martingale suitable weak solutions in hand (Section 4), we now focus on the second result of Theorem 2, that is, to establish the partial regularity of these martingale suitable weak solutions.
5.1 Blowing up analysis